OpenStudy (anonymous):

how do you do 3/5 to the negative 3

OpenStudy (anonymous):

\[(3/5)^-3\] = \[(5/3)^3\] =\[(5^3/3^3)\] =\[(125/9)\]

OpenStudy (anonymous):

you get it?

OpenStudy (anonymous):

yes thanks so much

OpenStudy (anonymous):

ok try to understand it, don't just copy answers. good luck